OpenStudy (anonymous):

y=x^2-2. What is Δ Y?

OpenStudy (ash2326):

Δ Y= \(dy\) let's find Δ y/Δx when Δ x--->0 Δ y/Δx ---> dy/dx let's find dy/dx =2x dy=2x dx so Δ y=2xΔx

OpenStudy (anonymous):

That wasn't one of the answers.. I chose something longer like [(x-Δx)^2-(something) (x^2-2)]

OpenStudy (ash2326):

yeah it can be like this also y=x^2-2 (y+Δ y)=(x+Δx)^2-2 subtracting the two , we get Δy= (x+Δx)^2-(x^2)

OpenStudy (anonymous):

Shoot, got that one wrong.

OpenStudy (ash2326):

are you writing a test?

OpenStudy (anonymous):

Nope, I just took my calc test and i'm checking the ones i wasn't sure about!

OpenStudy (ash2326):

okay :) do it well next time:)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Do you know if there is a function with 5 horizontal asymptotes? this was another t/f question

OpenStudy (ash2326):

We'd have to check options, can't tell like this . Maybe the function is piece wise defined